The famous WWII german gun know as the "88" was actually the bore size of the weapon in mm. Thus the "88" had an 88.0mm bore size. What is the calliber of this gun if the term caliber refers to the bore size in inches multiplied by 100(10.0mm=1.00cm; 2.54cm=1.00in)?
1
Expert's answer
2016-12-16T06:05:15-0500
Bore size in cm:

88.0 mm = 88.0 / 10 cm = 8.80 cm

Bore size in inches:

8.80 cm = 8.80 / 2.54 in = 3.465 in

Caliber:

3.465 in × 100 = 346.5